# On small perturbations of coherent information

### v1

@@ Abstract
Quantum capacity quantifies the amount of quantum information that can be transmitted by a quantum channel with an arbitrary small probability of error. Mathematically, the quantum capacity is given by an asymptotic formula involving the one-shot quantum capacity of the associated channel, which, due to purely quantum effects such as superadditivity of one-shot quantum capacity, is rarely computable. The one-shot quantum capacity is mathematically characterized in terms of optimizing an entropic quantity referred to as coherent information over all possible input states of a channel, the computation of which also tends to be intractable due to the difficulty of optimizing the coherent information. In this work, we develop perturbative methods for analyzing the behavior of coherent information of a quantum channel with respect to small perturbations of the input state. By doing so, we are able to derive three general criteria for determining whether an input state yields suboptimal coherent information. We then show how our criteria yield sufficient conditions for superadditivity of one-shot quantum capacity, and also for detecting a positive gap between one-shot private capacity and one-shot quantum capacity. The utility of our criteria is illustrated through examples, which yield new results regarding the one-shot quantum capacity of qubit depolarizing channels, Pauli channels and dephrasure channels.
@@ I Introduction
Classical information theory was established by the pioneering work of Shannon, who in a single paper laid the theoretical foundation for nearly all of modern communication technology [1]. A fundamental result of Shannon was the mathematical characterization of the maximum rate of information that can be transmitted by a classical noisy channel $\mathcal{N}$ with an arbitrarily low probability of error, which is referred to as the channel capacity of $\mathcal{N}$. If $X$ and $Y$ denote the input and output alphabets of a classical channel $\mathcal{N}$, and $\bm{p}(x)$ is a distribution on $X$, then the mutual information between sender and receiver is defined to be the non-negative real number $I(X:Y)$ given by
@@
where $\bm{p}(x,y)=\bm{p}(x)\bm{p}(y|x)$ is the associated joint distribution and $\bm{p}(y)$ is the induced distribution $Y$. What Shannon proved is that the channel capacity of $\mathcal{N}$ is the real number $C(\mathcal{N})$ given by
$C(\mathcal{N})=\max_{\bm{p}}\ I(X:Y)\,,$ (2)
where the maximum is taken over all input distributions $\bm{p}$ for the channel $\mathcal{N}$.
@@
where $Q^{(1)}(\mathcal{N})$ is the one-shot quantum capacity of a quantum channel $\mathcal{N}$, which represents the maximum rate of transmitting quantum information in a single use of the channel. The one-shot quantum capacity in turn is given by
@@
where $I_{c}(\rho,\mathcal{N})$ is the coherent information of $\mathcal{N}$ with respect to the input state $\rho$, which is a measure of correlation between the output system and the environment after using $\mathcal{N}$ to transmit $\rho$.
The formula for the one-shot quantum capacity (4) may be viewed as a quantum analog of formula (2) for the classical capacity, and as such, coherent information is often viewed as a quantum analog of the classical mutual information (1). However, while the classical capacity is additive, i.e., $C(\mathcal{N}_{1}\otimes\mathcal{N}_{2})=C(\mathcal{N}_{1})+C(\mathcal{N}_{2})$, the one-shot quantum capacity is superadditive, since for general quantum channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ we instead have the inequality
@@
Quantum channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ for which the above inequality (5) is strict exhibit holistic properties when combined in parallel which cannot be reduced to properties of its individual constituents, which is a characteristic feature of the transmission of quantum information. In particular, there are many known examples of quantum channels with superadditive one-shot quantum capacity [7, 11, 12, 13], and there even exists channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ with $Q(\mathcal{N}_{1})=Q(\mathcal{N}_{2})=0$ such that $Q(\mathcal{N}_{1}\otimes\mathcal{N}_{2})>0$, a phenomenon fist discovered by Smith and Yard which is now referred to as superactivation [14, 15]. Moreover, for every $n>0$, there exists a quantum channel $\mathcal{N}_{n}$ such that $Q^{(1)}(\mathcal{N}_{n}^{\otimes m})=0$ for all $m\leq n$, yet $Q(\mathcal{N}_{n})>0$ [16]. Such superadditive phenomena renders quantum capacity to be an elusive quantity which, apart from some special cases, is almost never explicitly computable. Not only this, but even calculating one-shot quantum capacity $Q^{(1)}$ also requires non-convex optimization over a continuum of states, which is usually an intractable problem.
@@
In this work, we further develop the perturbation methods employed in Refs. [17, 18] to derive general criteria for determining whether a quantum state yields suboptimal coherent information. As an application of our results, we show how our criteria yield sufficient conditions for detecting superadditivity of one-shot quantum capacity. To illustrate the utility of our criteria, we apply them to qubit depolarizing channels and qutrit platypus channels as studied in [19, 17, 20, 21], which turn out to yield some new results in such a context. For example, in the case of qubit depolarizing channels $\mathcal{D}_{p}$ with depolarizing parameter $p\in[0.2518,0.2524]$, it is shown in Refs. [22, 7, 23] that for all $n>2$ there exists states $\omega_{n}$ such that
$I_{c}(\omega_{n},\mathcal{D}_{p}^{\otimes n})>nQ^{(1)}(\mathcal{D}_{p})\,.$ (6)
However, it was unknown if there exists a state $\omega_{2}$ and some $p\in[0,1]$ for which inequality (6) holds for $n=2$. Applying our criteria, we produce such a state $\omega_{2}$ realizing the inequality (6) for $n=2$ in the range $0.146\leq p\leq 0.2524$, and moreover, our results extend the validity of inequality (6) to the same range.
@@ II Preliminaries
Given a quantum channel $\mathcal{N}$, there exists an environment $E$ and an isometric embedding $V:\mathcal{H}_{A}\to\mathcal{H}_{B}\otimes\mathcal{H}_{E}$ such that
@@
where $S(\cdot)$ denotes von Neumann entropy. It is straightforward to show that the coherent information is independent of the choice of an environment system $E$ which determines the complimentary channel $\mathcal{N}^{c}$, and thus is completely determined by the channel $\mathcal{N}$ and the input state $\rho$. The one-shot quantum capacity of $\mathcal{N}$ is then given by the formula
@@ III Perturbative analysis
As computing the one-shot quantum capacity $Q^{(1)}$ is generally intractable [26], in practice one often just selects some special input states which are believed to be nearly optimal for coherent information, yielding an approximation to one-shot quantum capacity for the subsequent information processing task. For example, for the mixed unitary channel $\mathcal{N}(\rho)=\sum_{i}p_{i}U_{i}\rho U_{i}^{\dagger}$, it is often the case that one simply chooses the maximally mixed state as its input, which is known to be optimal in certain examples. In this section, we fix a quantum channel $\mathcal{N}$ and an input state $\rho$, and we consider a perturbation of $\rho$ with respect to a small parameter $\varepsilon$. We then analyze the behavior of the associated coherent information under such perturbations, which enable us to derive three criteria for whether the state $\rho$ yields suboptimal coherent information. In particular, given a finite collection of states $\{\rho_{i}\}_{i=1}^{n}$ such that $\rho_{i}$ is optimal for a channel $\mathcal{N}_{i}$, our criteria may be used to determine if the product state $\rho=\rho_{1}\otimes\cdots\otimes\rho_{n}$ is not in fact optimal for the channel $\mathcal{N}=\mathcal{N}_{1}\otimes\cdots\otimes\mathcal{N}_{n}$.
@@
Suppose $\mathcal{N}(\rho)$ and $\mathcal{N}^{c}(\rho)$ are not both of not full rank, and suppose there exists a perturbative expansion $\rho(\varepsilon)$ of $\rho$ as in (8) such that
@@
The third criterion covers the case when $\mathcal{N}(\rho)$ and $\mathcal{N}^{c}(\rho)$ are both of full rank.
@@ IV Detecting superadditivity of one-shot quantum capacity
Proposition 2.
Let $\rho_{1}$ and $\rho_{2}$ be optimal states for quantum channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, let $\rho(\varepsilon)$ be a perturbative expansion of $\rho=\rho_{1}\otimes\rho_{2}$ as in (8), and let $f(\varepsilon)$ be the function given by (9) with $\mathcal{N}=\mathcal{N}_{1}\otimes\mathcal{N}_{2}$. Then a positive value of $f(\varepsilon)$ implies that the one-shot quantum capacity of $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ is super-additive, i.e., $Q^{(1)}(\mathcal{N}_{1}\otimes\mathcal{N}_{2})>Q^{(1)}(\mathcal{N}_{1})+Q^{(1)}(\mathcal{N}_{2})$.
@@
Example 1.
In this example we use Proposition 2 combined with Criterion 3 to establish results concerning superadditivity of one-shot quantum capacity of a qubit depolarizing channel, which for a depolarizing parameter $p\in[0,1]$ is given by
$\mathcal{D}_{p}(\omega)=(1-p)\omega+p\text{\rm Tr}(\omega)\mathds{1}/2\,.$ (15)
It is known that $Q^{(1)}(\mathcal{D}_{p})>0$ for $p\leq 0.2524$ [29], in which case the maximally mixed state $\rho=\mathds{1}/2$ is optimal. In Refs. [23, 7] it was established that in the range $0.2518\leq p\leq 0.255$, there exists states $\omega_{n}$ for all $n>2$ such that
$I_{c}(\omega_{n},\mathcal{D}_{p}^{\otimes n})>nQ^{(1)}(\mathcal{D}_{p})\,.$ (16)
For $n=2$ however it was unknown if there exists a state $\omega_{2}$ such that (16) holds for some $p\in[0,1]$. We now use Criterion 3 to show there indeed exists a state $\omega_{2}$ such that (16) holds for all $p\in[0.146,0.2524]$, thus implying superadditivity of one-shot quantum capacity for qubit depolarizing channels in the range $0.146\leq p\leq 0.2524$.
So let $q_{0}=(1-p)+p/4$ and $q_{k}=p/4$ for $k=1,2,3$. It is then straightforward to show that the complementary channel to $\mathcal{D}_{p}$ is given by $\mathcal{D}_{p}^{c}(\omega)=\text{\rm Tr}_{B}(V\omega V^{\dagger})$, where $V$ is the isometry given by $V=\sum_{k=0}^{3}q_{k}\sigma_{k}\otimes|k\rangle$ with $\sigma_{k}$ the $k$th Pauli matrix. Since $\mathcal{D}_{p}$ is unital, the commutation relations for Pauli matrices together with the fact that all $q_{k}$ are positive implies all traceless Hermitian matrices $A^{(1)}$ satisfy condition (13) of Criterion 3. We now show that condition (14) is also satisfied.
For this, let $\mathcal{N}=\mathcal{D}_{p}\otimes\mathcal{D}_{p}$. Since the maximally mixed state is optimal for $\mathcal{D}_{p}$ for $p\leq 0.2524$, we let $\rho=\mathds{1}/4$. We then consider the perturbative expansion $\rho(\varepsilon)=\rho+\varepsilon A^{(1)}$, where $A^{(1)}=|\psi\rangle\langle\psi|-\rho$ with $|\psi\rangle=(|00\rangle+|10\rangle)/\sqrt{2}$. By direct calculation we find
$\text{\rm Tr}(W_{\mathcal{N}})=[1+(1-p)^{2}]^{2}-1\,,$
and
$\text{\rm Tr}(W_{\mathcal{N}^{c}})$ $=q_{0}^{2}+12q_{0}q_{1}+25q_{1}^{2}$
$+\frac{4q_{0}q_{1}(q_{0}^{2}+7q_{0}q_{1}+5q_{1}^{2})}{(1-p)(q_{0}+q_{1})}\log\frac{q_{0}}{q_{1}}\,.$
[…]
@@
with $20<s\leq 1/$ 2. The qutrit platypus channel has been studied extensively in Refs. [20, 21], where it was shown that a state of the form $\rho(p)=(1-p)|0\rangle\langle 0|+p|2\rangle\langle 2|$ is optimal with respect to $\mathcal{N}_{s}$, so that
@@
As the operator $A^{(1)}$ given here satisfies condition (11) of Criterion 2, we now show the condition (12) also holds: Let $\mathcal{N}=\mathcal{N}_{s}\otimes\mathcal{A}_{\gamma}$, so that its complement is $\mathcal{N}^{c}=\mathcal{N}_{s}^{c}\otimes\mathcal{A}_{\gamma}^{c}$. We then have
@@
Since $a^{2}$ can be chosen to close to $1$ for $\varepsilon$ close to 00, it follows that condition (12) of Criterion 2 holds for all $\gamma\in[\frac{1}{2},\frac{1-s+sw}{2-2s-w+2sw})$, which together with Proposition 2 implies that the one-shot quantum capacity of $\mathcal{N}_{s}$ and $\mathcal{A}_{\gamma}$ is superadditive. We note that the perturbative expansion $\rho(\varepsilon)$ given by (17) is different from the perturbative state $\sigma(\varepsilon)$ utilized in Refs. [17, 21] to establish superadditivity of $Q^{(1)}$, which is given by
@@
and is thus rarely computable due not only to superadditive effects, but also due to intractable optimization procedures needed for the calculation of $\mathcal{P}^{(1)}$. However, it is still possible to detect a positive gap between $P^{(1)}$ and $Q^{(1)}$ without explicitly computing their exact values. In particular, we now show how our criteria implies a sufficient condition for $P^{(1)}$ to be strictly larger than $Q^{(1)}$:
@@
Now let $\rho$ be a pure state, and let $r>0$ be such that $\sigma-r\cdot\rho(\varepsilon)$ is positive semi-definite for sufficiently small $\varepsilon$ (if the kernel of $\sigma$ is contained in the kernel of $\rho(\varepsilon)$ for some $\varepsilon$, then such an $r$ necessarily exists). Now let $\{q_{i},|\psi_{i}\rangle\langle\psi_{i}|\}$ be an ensemble decomposition of $\sigma-r\cdot\rho(\varepsilon)$ corresponding to its spectral decomposition with respect to 1-dimensional projectors (pure states), so that $\{(r,\rho(\varepsilon)),(q_{i},|\psi_{i}\rangle\langle\psi_{i}|)\}$ is an ensemble decomposition of $\sigma$. We then have
@@
Let $\mathcal{N}=\mathcal{D}_{p}$, where $\mathcal{D}_{p}$ is the qubit depolarizing channel as given by (15). It is known that the maximally mixed state $\sigma=\mathds{1}/2$ is optimal for $\mathcal{N}$ for all $p\in[0,0.2524]$. Now consider the perturbative state $\rho(\varepsilon)$ given by
@@
[…]
In particular, since the complementary channel $\mathcal{N}_{p,q}^{c}$ has positive one-shot quantum capacity for all $(p,q)\in\mathcal{R}_{1}$, it follows that $P^{(1)}(\mathcal{N}_{p,q})-Q^{(1)}(\mathcal{N}_{p,q})>0$ for all $(p,q)\in\mathcal{R}_{1}$. This extends a result in Ref. [25], where it was shown such a positive gap exists in the range $p\in[0.08,0.125]$ and $q=3p$.
@@
thus the gap between $P^{(1)}$ and $C^{(1)}$ is at least $\chi(\{p_{i},\rho_{i}\},\mathcal{N}^{c})\geq 0$. According to the equality condition for concavity of entropy, it follows that $\chi(\{p_{i},\rho_{i}\},\mathcal{N}^{c})=0$ if and only if $\rho_{i}=\rho$ for all $i$, which implies that $P^{(1)}(\mathcal{N})=Q^{(1)}(\mathcal{N})=0$. Therefore, there exists a positive gap between $C^{(1)}(\mathcal{N})$ and $P^{(1)}(\mathcal{N})$ if $P^{(1)}(\mathcal{N})>Q^{(1)}(\mathcal{N})$, and this condition can be determined by Proposition 4.

### v2

@@ Abstract
Quantum capacity quantifies the amount of quantum information that can be transmitted by a quantum channel with an arbitrary small probability of error. Mathematically, the quantum capacity is given by an asymptotic formula involving the one-shot quantum capacity of the associated channel, which, due to purely quantum effects such as superadditivity of one-shot quantum capacity, is rarely computable. The one-shot quantum capacity is mathematically characterized in terms of optimizing an entropic quantity referred to as coherent information over all possible input states of a channel, the computation of which also tends to be intractable due to the difficulty of optimizing the coherent information. In this work, we develop perturbative methods for analyzing the behavior of first and second order derivatives of coherent information of a quantum channel with respect to small perturbations of the input state. By doing so, we are able to derive three general criteria for determining whether an input state yields suboptimal coherent information. We then show how our criteria yield sufficient conditions for superadditivity of one-shot quantum capacity, and also for detecting a positive gap between one-shot private capacity and one-shot quantum capacity. The utility of our criteria is illustrated through examples, which yield new results regarding the one-shot quantum capacity of qubit depolarizing channels, Pauli channels and dephrasure channels.
@@ I Introduction
Classical information theory was established by the pioneering work of Shannon, who in a single paper laid the theoretical foundation for nearly all of modern communication technology [1]. A fundamental result of Shannon was the mathematical characterization of the maximum rate of information that can be transmitted by a classical noisy channel $N$ with an arbitrarily low probability of error, which is referred to as the channel capacity of $N$. If $X$ and $Y$ denote the input and output alphabets of a classical channel $N$, and $\bm{p}(x)$ is a distribution on $X$, then the mutual information between sender and receiver is defined to be the non-negative real number $I(X:Y)$ given by
@@
where $\bm{p}(x,y)=\bm{p}(x)\bm{p}(y|x)$ is the associated joint distribution and $\bm{p}(y)$ is the induced distribution $Y$. What Shannon proved is that the channel capacity of $N$ is the real number $C_{\mathrm{Shan}}(N)$ given by
$C_{\mathrm{Shan}}(N)=\max_{\bm{p}}\ I(X:Y)\,$, (2)
where the maximum is taken over all input distributions $\bm{p}$ for the channel $N$.
@@
where $Q^{(1)}(\mathcal{N})$ is the one-shot quantum capacity of a quantum channel $\mathcal{N}$, which is given by
@@
where $I_{c}(\rho,\mathcal{N})$ is the coherent information of $\mathcal{N}$ with respect to the input state $\rho$. The coherent information $I_{c}(\rho,\mathcal{N})$ is a measure of correlation between the output system and the environment after using $\mathcal{N}$ to transmit $\rho$, and is a fundamental measure of quantum information [11].
The formula for the one-shot quantum capacity (4) may be viewed as a quantum analog of formula (2) for the classical capacity, and as such, coherent information is often viewed as a quantum analog of the classical mutual information (1). However, while the classical capacity is additive, i.e., $C_{\mathrm{Shan}}(N_{1}\times N_{2})=C_{\mathrm{Shan}}(N_{1})+C_{\mathrm{Shan}}(N_{2})$, the one-shot quantum capacity is superadditive, since for general quantum channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ we instead have the inequality
@@
Quantum channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ for which the above inequality (5) is strict exhibit holistic properties when combined in parallel which cannot be reduced to properties of its individual constituents, which is a characteristic feature of the transmission of quantum information. In particular, there are many known examples of quantum channels with superadditive one-shot quantum capacity [7, 12, 13, 14], and there even exists channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ with $Q(\mathcal{N}_{1})=Q(\mathcal{N}_{2})=0$ such that $Q(\mathcal{N}_{1}\otimes\mathcal{N}_{2})>0$, a phenomenon first discovered by Smith and Yard which is now referred to as superactivation [15, 16]. Moreover, for every $n>0$, there exists a quantum channel $\mathcal{N}_{n}$ such that $Q^{(1)}(\mathcal{N}_{n}^{\otimes m})=0$ for all $m\leq n$, yet $Q(\mathcal{N}_{n})>0$ [17]. In addition, superadditivity appears in many information transmission scenarios using quantum channels, indicating that the corresponding channel capacity is non-additive in general. For instance, there are examples showing that superadditivity exists in one-shot private capacity $P^{(1)}$ [18, 19], and private capacity $P$ [20, 21]. Furthermore, such superadditive phenomena renders quantum channel capacity to be an elusive quantity which, apart from some special cases, is almost never explicitly computable. Not only this, but even calculating one-shot capacity also requires non-convex optimization over a continuum of states, which is usually an intractable problem.
Although superadditivity generally holds, it’s still possible to find special quantum channels $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ satisfying additivity: $Q^{(1)}(\mathcal{B}_{1}\otimes\mathcal{B}_{2})=Q^{(1)}(\mathcal{B}_{1})+Q^{(1)}(\mathcal{B}_{2})$, which means that we can bypass regularized expressions and obtain the channel capacity by calculating the one-shot capacity of the channel (called single letter formula). For example, in scenarios where quantum channels are used to transmit quantum information, the degradable channel [9, 22] has been shown with additive coherent information when combined with another degradable channels, and only entanglement-breaking channels are known with additive coherent information and quantum capacity combined with any other channels [23].
Recently, Refs. [24, 25] propose sufficient conditions for determining the positivity of one-shot quantum capacity for generic quantum channels. Given a general quantum channel $\mathcal{N}$, their fundamental idea involves perturbing a pure state to obtain a quantum state $\rho(\varepsilon)$, then considering the derivative of the coherent information for the channel $\mathcal{N}$ with respect to the state $\rho(\varepsilon)$, $I_{c}^{\prime}(\varepsilon):=\frac{\mathrm{d}I_{c}[\mathcal{N}(\rho(\varepsilon))]}{\mathrm{d}\varepsilon}$. Since the coherent information of the general channel with respect to any pure state is zero, if this derivative is greater than zero, it proves that there exists a quantum state $\rho(\varepsilon)$ such that the coherent information $I_{c}(\rho,\mathcal{N})$ is positive implying positivity of (one-shot) quantum capacity of the channel $\mathcal{N}$.
In this work, we further develop the perturbation methods employed in Refs. [24, 25] to derive general criteria for determining whether a quantum state yields suboptimal coherent information. Specifically, given a generic channel $\mathcal{N}$ and a quantum state $\rho$ (which may be mixed), we consider the perturbative expansion $\rho(\varepsilon)$ given in Eq. (7) below, then compute the first and second orders derivative of the function $f(\varepsilon)$ which is defined as the difference in coherent information between the quantum state before and after perturbation. It’s clear that the positivity of $f(\varepsilon)$ implies that there exists a quantum state $\rho(\varepsilon)$ with larger coherent information than $\rho$. Based on this observation, we provide general criteria for determining whether there exists a positive value of $f(\varepsilon)$ according to the positivity of its first and second order derivatives.
@@
As an application of our results, we show how our criteria yield sufficient conditions for detecting superadditivity of one-shot quantum capacity. To illustrate the utility of our criteria, we apply them to qutrit platypus channels as studied in [26, 24, 27, 23], which turn out to yield some new results in such a context. For example, in the case of qutrit platypus channels combined with amplitude damping channels, we provide various input state forms in Eq. (14), which realizes superadditive one-shot quantum capacity by leveraging Criterion 2.
@@ II Preliminaries
Given a quantum channel $\mathcal{N}$, there exists an environment $E$ and an isometry $V:\mathcal{H}_{A}\to\mathcal{H}_{B}\otimes\mathcal{H}_{E}$ such that
@@
where $S(\rho)=-\text{\rm Tr}(\rho\log\rho)$ denotes von Neumann entropy. It is straightforward to show that the coherent information is independent of the choice of an environment system $E$ which determines the complimentary channel $\mathcal{N}^{c}$, and thus is completely determined by the channel $\mathcal{N}$ and the input state $\rho$. The one-shot quantum capacity of $\mathcal{N}$ is then given by the formula
@@ III Perturbative analysis
As computing the one-shot quantum capacity $Q^{(1)}$ is generally intractable [30], in practice one often selects some special input states which are believed to be nearly optimal for coherent information, yielding an approximation to one-shot quantum capacity for the subsequent information processing task. For example, for a mixed unitary channel $\mathcal{N}(\rho)=\sum_{i}p_{i}U_{i}\rho U_{i}^{\dagger}$, it is often the case that one simply chooses the maximally mixed state as its input, which is known to be optimal in certain examples. In this section, we fix a quantum channel $\mathcal{N}$ and an input state $\rho$, and we consider a perturbation of $\rho$ with respect to a small parameter $\varepsilon$. We then analyze the behavior of the associated coherent information under such perturbations, which enable us to derive three criteria for whether the state $\rho$ yields suboptimal coherent information. In particular, given a finite collection of states $\{\rho_{i}\}_{i=1}^{n}$ such that $\rho_{i}$ is optimal for a channel $\mathcal{N}_{i}$, our criteria may be used to determine if the product state $\rho=\rho_{1}\otimes\cdots\otimes\rho_{n}$ is not in fact optimal for the channel $\mathcal{N}=\mathcal{N}_{1}\otimes\cdots\otimes\mathcal{N}_{n}$.
@@
Suppose $\mathcal{N}(\rho)$ and $\mathcal{N}^{c}(\rho)$ are not both of full rank, and suppose there exists a perturbative expansion $\rho(\varepsilon)$ of $\rho$ as in (7) such that
@@
The third criterion covers the case when $\mathcal{N}(\rho)$ and $\mathcal{N}^{c}(\rho)$ are both of full rank. Since whether these two operators are full rank depends on the dimension of the corresponding space, we must emphasize that the dimensions of output $d^{*}_{\text{out}}(\mathcal{N})$ and environment $d^{*}_{\text{env}}$ system of the channel $\mathcal{N}$ are both minimal, in the sense that $d^{*}_{\text{out}}(\mathcal{N})=\mathrm{rank}[\mathcal{N}(\mathds{1})]$ and $d^{*}_{\text{env}}=\mathrm{rank}[\mathcal{N}^{c}(\mathds{1})]$ as shown in Lemma II.3 of [25], and hence the ranks of $\mathcal{N}(\rho)$ and $\mathcal{N}^{c}(\rho)$ satisfy $\mathrm{rank}[\mathcal{N}(\rho)]=\mathrm{rank}[\mathcal{N}(\mathds{1})]$ and $\mathrm{rank}[\mathcal{N}^{c}(\rho)]=\mathrm{rank}[\mathcal{N}^{c}(\mathds{1})]$.
@@
When $\mathcal{N}(\rho)$ and $\mathcal{N}^{c}(\rho)$ are both of full rank, Ref. [33] computes the gradient and Hessian of ($n$-shot) coherent information of the strictly positive channel $\mathcal{N}$ at a positive definite state $\rho$. In contrast, our results propose explicit criteria for determining whether a quantum state is optimal for coherent information.
@@ IV Detecting superadditivity of one-shot quantum capacity
Proposition 2 ([24]).
Let $\rho_{1}$ and $\rho_{2}$ be optimal states for quantum channels $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, let $\rho(\varepsilon)$ be a perturbative expansion of $\rho=\rho_{1}\otimes\rho_{2}$ as in (7), and let $f(\varepsilon)$ be the function given by (8) with $\mathcal{N}=\mathcal{N}_{1}\otimes\mathcal{N}_{2}$. Then a positive value of $f(\varepsilon)$ implies that the one-shot quantum capacity of $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ is superadditive, i.e., $Q^{(1)}(\mathcal{N}_{1}\otimes\mathcal{N}_{2})>Q^{(1)}(\mathcal{N}_{1})+Q^{(1)}(\mathcal{N}_{2})$.
@@
with $0<s\leq 1/2$. The qutrit platypus channel has been studied extensively in Refs. [27, 23], where it was shown that a state of the form $\rho(p)=(1-p)|0\rangle\langle 0|+p|2\rangle\langle 2|$ is optimal with respect to $\mathcal{N}_{s}$, so that
@@
As the output states $\mathcal{N}_{1}(\rho_{1})\otimes\mathcal{N}_{2}(\rho_{2})$ and $\mathcal{N}_{1}^{c}(\rho_{1})\otimes\mathcal{N}_{2}^{c}(\rho_{2})$ are both diagonal which implies that the projectors $P_{i}$ and $Q_{j}$ in condition (10) of Criterion 2 are corresponding to the computational basis, and the diagonal entry of operators $\mathcal{N}_{1}\otimes\mathcal{N}_{2}(A^{(1)})$ and $\mathcal{N}_{1}\otimes\mathcal{N}_{2}(A^{(1)})$ are all zero, one can know that the operator $A^{(1)}$ given here satisfies condition (10) of Criterion 2, we now show the condition (11) also holds: Let $\mathcal{N}=\mathcal{N}_{s}\otimes\mathcal{A}_{\gamma}$, so that its complement is $\mathcal{N}^{c}=\mathcal{N}_{s}^{c}\otimes\mathcal{A}_{\gamma}^{c}$. We then have
@@
Since $a^{2}$ can be chosen to close to $1$ for $\varepsilon$ close to 0, it follows that condition (11) of Criterion 2 holds for all $\gamma\in[\frac{1}{2},\frac{1-s+sw}{2-2s-w+2sw})$, which is the same range of $\gamma$ provided in [24, 23], together with Proposition 2 implies that the one-shot quantum capacity of $\mathcal{N}_{s}$ and $\mathcal{A}_{\gamma}$ is superadditive. We note that the perturbative expansion $\rho(\varepsilon)$ given by (14) is different from the perturbative state $\sigma(\varepsilon)$ utilized in Refs. [24, 23] to establish superadditivity of $Q^{(1)}$, which is given by
@@
As for the applicability of Criterion 3, we have yet to find any examples of pairs $(\rho,\mathcal{N})$ for which the conditions (12) and (13) of Criterion 3 are both satisfied. It could be the case that such conditions are satisfied only on a set of measure-zero, in which case the utility of Criterion 3 would be quite limited. As such, it would be interesting to determine the locus of pairs $(\rho,\mathcal{N})$ satisfying the conditions (12) and (13) for fixed input and output systems of the channel $\mathcal{N}$.
@@
and is thus rarely computable due not only to superadditive effects, but also due to intractable optimization procedures needed for the calculation of $\mathcal{P}^{(1)}$. However, it is still possible to detect a positive gap between $P^{(1)}$ and $Q^{(1)}$ without explicitly computing their exact values. While Watanabe [28] proved that $Q^{(1)}(\mathcal{B})=P^{(1)}(\mathcal{B})$ if the one-shot quantum capacity of the complementary channel $\mathcal{B}^{c}$ vanishes [37, 38]. There exist quantum channels displaying the positive gap between one-shot quantum and private capacity, such as Horodecki channels [39, 40, 41, 42], half-rocket channels [43], platypus channels [26, 44, 27, 45] and dephrasure channels [29]. In particular, we now show how our criteria implies a sufficient condition for $P^{(1)}$ to be strictly larger than $Q^{(1)}$ which is more general than the result of Watanabe [28]:
@@
Now let $\rho$ be a pure state, and let $r>0$ be such that $\sigma-r\cdot\rho(\varepsilon)$ is positive semi-definite for sufficiently small $\varepsilon$ (if the kernel of $\sigma$ is contained in the kernel of $\rho(\varepsilon)$ for some $\varepsilon$, then such an $r$ necessarily exists). Now let $\{q_{i},|\psi_{i}\rangle\langle\psi_{i}|\}$ be the spectral decomposition of $\sigma-r\cdot\rho(\varepsilon)$ with respect to 1-dimensional projectors (pure states), so that $\{(r,\rho(\varepsilon)),(q_{i},|\psi_{i}\rangle\langle\psi_{i}|)\}$ is an ensemble decomposition of $\sigma$. We then have
@@
Let $\mathcal{N}=\mathcal{D}_{p}$, where $\mathcal{D}_{p}$ is the qubit depolarizing channel $\mathcal{D}_{p}(\rho)=(1-p)\rho+p\,\mathds{1}/2$. It is known that the maximally mixed state $\sigma=\mathds{1}/2$ is optimal for $\mathcal{N}$ for all $p\in[0,0.2524]$. Now consider the perturbative state $\rho(\varepsilon)$ given by
@@
$\mathcal{R}_{1}=\left\{(p,q)\,\,\Big|\,\,p\in[0,1/2],\,\,0\leq q<\frac{(1-2p)^{2}}{1+(1-2p)^{2}}\right\}\,.$
In particular, when $(p,q)\in\mathcal{R}_{1}$, the complementary channel $\mathcal{N}_{p,q}^{c}$ has positive one-shot quantum capacity, and the optimal state for $\mathcal{N}_{p,q}$ is of full rank. Therefore, for any state $\rho$ such that $I_{c}(\rho,\mathcal{N}^{c}_{p,q})>0$, we can always find a small parameter $r$ such that $\sigma-r\cdot\rho>0$. Then it follows that $P^{(1)}(\mathcal{N}_{p,q})-Q^{(1)}(\mathcal{N}_{p,q})>0$ for all $(p,q)\in\mathcal{R}_{1}$. This extends a result in Ref. [29], where it was shown such a positive gap exists in the range $p\in[0.08,0.125]$ and $q=3p$.
@@
thus the gap between $P^{(1)}$ and $C^{(1)}$ is at least $\chi(\{p_{i},\rho_{i}\},\mathcal{N}^{c})\geq 0$. According to the equality condition for concavity of entropy, it follows that $\chi(\{p_{i},\rho_{i}\},\mathcal{N}^{c})=0$ if and only if $\mathcal{N}^{c}(\rho_{i})=\mathcal{N}^{c}(\rho)$ for all $i$. Consequently, the perturbative method may be of limited utility in identifying the gap between the one-shot private capacity and the Holevo capacity, analogous to its failure in detecting the nonadditivity of the minimum output entropy [46].